# Evaluating Retrieval-Augmented Generation Systems on Unanswerable, Uncheatable, Realistic, Multi-hop Queries

### v2

@@ Abstract
Real-world use cases often present RAG systems with complex queries for which relevant information is missing from the corpus or is incomplete. In these settings, RAG systems must be able to reject unanswerable, out-of-scope queries and identify failures of retrieval and multi-hop reasoning. Despite this, existing RAG benchmarks rarely reflect realistic task complexity for multi-hop or out-of-scope questions, which often can be cheated via disconnected reasoning (i.e., solved without genuine multi-hop inference) or require only simple factual recall. This limits the ability for such benchmarks to uncover limitations of existing RAG systems. To address this gap, we present the first pipeline for automatic, difficulty-controlled creation of uncheatable, realistic, unansw-erable, and multi-hop queries (CRUMQs), adaptable to any corpus and domain. We use our pipeline to create CRUMQs over two popular RAG datasets and demonstrate its effectiveness via benchmark experiments on leading retrieval-augmented LLMs. Results show that compared to prior RAG benchmarks, CRUMQs are highly challenging for RAG systems and achieve up to 81.0% reduction in cheatability scores. More broadly, our pipeline offers a simple way to enhance benchmark difficulty and realism and drive development of more capable RAG systems.
@@ 1 Introduction
Retrieval Augmented Generation (RAG) [23, 11] is a powerful approach for many NLP tasks, enabling LLMs to respond to diverse user requests by leveraging an external document collection. While RAG is highly effective at increasing model credibility [15], mitigating hallucinations, and improving response quality [12], there remains a need to better understand how such systems handle complex, multi-part queries when available information from the corpus is insufficient. In particular, RAG systems must be able to appropriately reject unanswerable queries and localize retrieval or reasoning failures when responding to multi-hop requests [37, 29]. These capabilities are crucial for reliable deployment of RAG systems in high-stakes domains where information is often missing or incomplete.
Existing RAG benchmarks [44, 14] rarely evaluate systems’ ability to handle realistic multi-hop or unanswerable queries. Multi-hop RAG benchmarks [37, 21] focus on synthetic task domains or suffer from disconnected reasoning [38, 39] wherein shortcuts can be exploited to achieve correct answers, while unanswerable RAG benchmarks [29] remain limited to simplistic factual recall settings.
@@ 3 Method
We follow the pipeline shown in Fig. 1 to generate CRUMQs. We begin by extracting relevant topic keyphrases over the provided document collection. In this work, we assume the collection is a RAG corpus $D$ with associated information-seeking requests and gold retrieved documents from $D$ per request. However, our pipeline may be generalized by using synthetic requests or topic modeling.
@@
Step I. Topics are extracted via two simple steps. A frontier LLM is few-shot prompted to extract short keyphrases from each information-seeking request. [^1] Our code and prompts are available at https://github.com/pybeebee/CRUMQs. To obtain finer-grained topics, we then pair each initial topic with a gold document for the request and employ a second few-shot prompt to extract document-grounded topics. To ensure topics are sufficiently distinct and have good coverage of the corpus, we perform a deduplication step via embedding similarity. We use the BAAI/bge-large-en-v1.5 embedder with a similarity threshold of 0.95.
@@
Step III. The externally sourced articles are then used to generate queries that are either fully unanswerable (relevant information is not found in $D$, only in the externally sourced documents) or partially unanswerable (relevant information is in $D$, but at least one key fact or detail required to provide a complete and correct answer is absent). To do so, we first split each gold article and each external article into 1,024-token chunks via LangChain [8], tracking for each chunk its source, URL, associated topic keyphrase, and associated request. Chunks are filtered for relevance to the topic and request via a binary LLM judgment. We then construct the set of all possible groups of 2-6 chunks such that at least one externally sourced chunk is in each group. As this set may be overly large in practice, we place a limit $N_{c}$ on the number of contexts that are created for each total number of chunks and each ratio of external:gold chunks. To obtain seed queries, we prompt a strong generator LLM to systematically create up to 10 multi-hop QA pairs using each multi-chunk context, [^2] We found that QA generation with multi-chunk contexts was more fruitful than with extracted claims [37] or entity-relation triplets [22] from document chunks. making sure information is leveraged across chunk boundaries. For fully unanswerable queries we require all chunks in a context to be externally sourced; for partially unanswerable queries we require at least one gold chunk. We ensure diverse task formulations by adapting prompts from prior work in multi-document QA generation [25, 42].
@@
Step V. The data is finally filtered to ensure only truly multi-hop and high-quality queries remain. Following [24], we first annotate each QA pair with a synthetic chain-of-thought (CoT) explanation of the answer in the oracle setting (i.e., assuming both gold and externally sourced documents are available), and record the number of hops required to solve the question. Queries that do not adhere to the intended hop count are discarded. [^3] We retain a selection of single-hop queries in §4 for comparability to prior works. We then utilize a strong LLM to assess each unanswerable QA pair in the oracle setting according to the following criteria [13, 37]: context necessity, context sufficiency, answer correctness, answer uniqueness. Contextual criteria are additionally assessed assuming only the given corpus is available. We utilize the same scoring scale (Likert, 0-2) as [13]: only QA pairs which receive a score of at least 1 for all 6 criteria are retained. [^4] We manually review a subset of examples to validate the LLM verification results.
@@ 4 Experimental Setup
Datasets. Using our pipeline, we first generate CRUMQs over NeuCLIR [6] and TREC RAG 2025 [9], two popular RAG datasets which each consist of a large document collection with associated user requests. We focus on English texts, set $N_{e}=200$ and $N_{c}=50$, and use Llama3.3-70B-Instruct [18] as the generation and verification model to balance costs and quality. [^5] As we show in §5, creating effective CRUMQs does depend on proprietary models. This leads to a total of 3,048 CRUMQs. We next use UAEval4RAG to generate baseline out-of-database RAG queries over NeuCLIR and TREC RAG 2025, leading to 7,559 out-of-database queries. Finally, we utilize the MultiHop-RAG dataset as-is, which consists of 2,556 multi-hop RAG queries and their associated gold contexts.
Unanswerability Evaluation. To verify the unanswerability value of our pipeline, we adopt the same experimental setup as in [29] to benchmark four leading RAG systems on the CRUMQs and UAEval4RAG queries, downsampled to 3,048 examples for fair comparison. We assume no access to external documents and use LlamaIndex [27] for RAG system implementation. We additionally evaluate the performance of RAG systems that use leading proprietary LLMs Gemini-2.5-Pro [17] and GPT-5 for generation, in addition to GPT-4o as in [29].
Cheatability Evaluation. To demonstrate the robustness of our queries against reasoning shortcuts, we compare our CRUMQs (downsampled to 2,556 examples) against MultiHop-RAG [37]. Following [38], we first obtain LLM predictions for each dataset in the oracle setting on the tasks of answer prediction and paragraph-level support identification. We use Llama3.1-8B-Instruct, Llama3.3-70B-Instruct, GPT-4o, and Gemini-2.5-Pro to ensure coverage of diverse model types and sizes. We then repeat the same experiment on the DiRe probe [38] of each dataset, which is a model-agnostic probe to gauge the extent of disconnected reasoning; additional details are in [38].
@@ 5 Results and Analysis
Table 1: Performance of RAG systems on CRUMQs vs. UAEval4RAG queries, using GPT-4o for generation. denotes Holm-Bonferroni corrected significance ($p<0.05$).
@@
Table 2: Performance of proprietary model RAG systems on CRUMQs vs. UAEval4RAG queries. As in [29], no reranking or rewriting is used. denotes Holm-Bonferroni corrected significance ($p<0.05$).
@@
Compared to prior multi-hop RAG benchmarks, CRUMQs are significantly more difficult and much less cheatable, with all comparisons significant after Holm–Bonferroni correction ($p<0.05$). As indicated by the black points in Fig. 3, we observe that CRUMQs pose a much greater challenge for leading LLMs than MultiHop-RAG across all tasks, with models achieving answer prediction scores up to only 31.9 versus up to 84.4 for MultiHop-RAG. Moreover, up to 96.7% of queries in MultiHop-RAG can be answered via disconnected reasoning, compared to only up to 23.6% of our CRUMQs. A similar trend is observed for the support identification task. These results confirm the efficacy of our pipeline for creation of challenging multi-hop RAG queries.

### v3

# Investigating Retrieval-Augmented Generation Systems on Unanswerable, Uncheatable, Realistic, Multi-hop Queries

@@ Abstract
Real-world use cases often present RAG systems with complex queries for which relevant information is missing from the corpus or is incomplete. In these settings, RAG systems must be able to reject unanswerable, out-of-scope queries and identify failures of retrieval and multi-hop reasoning. Despite this, existing RAG benchmarks rarely reflect realistic task complexity for multi-hop or out-of-scope questions, which often can be cheated via disconnected reasoning (i.e., solved without genuine multi-hop inference) or require only simple factual recall. This limits the ability for such benchmarks to uncover limitations of existing RAG systems. To address this gap, we present the first pipeline for automatic, difficulty-controlled creation of uncheatable, realistic, unansw-erable, and multi-hop queries (CRUMQs), adaptable to any corpus and domain. We use our pipeline to create CRUMQs over two popular RAG datasets and demonstrate its effectiveness via benchmark experiments on leading retrieval-augmented LLMs. Results show that compared to prior RAG benchmarks, CRUMQs are highly challenging for RAG systems and achieve up to 81.0% reduction in cheatability scores. More broadly, our pipeline offers a simple way to enhance benchmark difficulty and drive development of more capable RAG systems.
@@ 1 Introduction
Retrieval Augmented Generation (RAG) [23, 11] is a powerful approach for many NLP tasks, enabling LLMs to respond to diverse user requests by leveraging an external document collection. While RAG is highly effective at increasing model credibility [15], mitigating hallucinations, and improving response quality [12], there remains a need to better understand how such systems handle complex, multi-part queries when available information from the corpus is insufficient. In particular, RAG systems must be able to appropriately reject unanswerable queries (i.e., those for which no relevant information is present in the corpus) and localize retrieval or reasoning failures when responding to multi-hop requests [37, 29]. These capabilities are crucial for reliable deployment of RAG systems in high-stakes domains where information is often missing or incomplete.
Existing RAG benchmarks [44, 14] rarely evaluate systems’ ability to handle realistic multi-hop or unanswerable queries. Multi-hop RAG benchmarks [37, 21] focus on synthetic tasks or can be cheated via disconnected reasoning [38, 39] wherein shortcuts can be exploited to achieve correct answers, while unanswerable RAG benchmarks [29] remain limited to simplistic factual recall settings.
@@ 3 Method
We follow the pipeline shown in Fig. 1 to generate CRUMQs. We begin by extracting relevant topic keyphrases over the provided document collection. In this work, we assume the collection is a RAG corpus $D$ with associated information-seeking requests and gold relevant documents from $D$ per request. However, our pipeline may be generalized by using synthetic requests or topic modeling.
@@
Step I. Topics are extracted via two simple steps. A frontier LLM is few-shot prompted to extract short keyphrases from each information-seeking request. [^1] Our code and prompts are available at https://github.com/pybeebee/CRUMQs. To obtain finer-grained topics, we then pair each initial topic with a gold document for the request and employ a second few-shot prompt to extract document-grounded topics. To ensure topics are sufficiently distinct and have good coverage of the corpus, we perform a deduplication step via embedding similarity. [^2] Other similarly effective deduplication methods may be used, such as a LLM call. We use the BAAI/bge-large-en-v1.5 embedder with a similarity threshold of 0.95.
@@
Step III. The externally sourced articles are then used to generate queries that are either fully unanswerable (relevant information is not found in $D$, only in the externally sourced documents) or partially unanswerable (relevant information is in $D$, but at least one key fact or detail required to provide a complete and correct answer is absent).
To do so, we first split each gold article and each external article into 1,024-token chunks via LangChain [8], tracking for each chunk its source, URL, associated topic keyphrase, and associated request. Chunks are filtered for relevance to the topic and request via a binary LLM judgment. We then construct the set of all possible groups of 2-6 chunks such that between 1-5 externally sourced chunks are in each group. As this set may be overly large in practice, we place a limit $N_{c}$ on the number of contexts that are created for each total number of chunks and each ratio of external:gold chunks.
To obtain seed queries, we prompt a strong generator LLM to systematically create up to 10 multi-hop QA pairs using each multi-chunk context, [^3] While some works explore QA generation over extracted claims [37] or entity-relation triplets [22] from chunks, we found direct use of chunks to be more fruitful. making sure information is leveraged across chunk boundaries. For fully unanswerable queries we require all chunks in a context to be externally sourced; for partially unanswerable queries we require at least one gold chunk. Systematicity and diversity of resulting task formulations are ensured by adapting prompts from prior multi-document QA pipelines [25, 42].
@@
Step V. The data is finally filtered to ensure only truly multi-hop and high-quality queries remain. Following [24], we first annotate each QA pair with a synthetic chain-of-thought (CoT) explanation of the answer in the oracle setting (i.e., assuming both gold and externally sourced documents are available), and record the number of hops required to solve the question. Queries that do not adhere to the intended hop count are discarded. [^4] We retain a selection of single-hop queries in §4 for comparability to prior works. We then utilize a strong LLM to assess each unanswerable QA pair in the oracle setting according to the following criteria [13, 37]: context necessity, context sufficiency, answer correctness, answer uniqueness. Contextual criteria are additionally assessed assuming only the given corpus is available. We utilize the same scoring scale (Likert, 0-2) as [13]: only QA pairs which receive a score of at least 1 for all 6 criteria are retained. A subset of examples is manually reviewed to validate the LLM verification results.
@@ 4 Experimental Setup
Datasets. Using our pipeline, we first generate CRUMQs over NeuCLIR [6] and TREC RAG 2025 [9], two popular RAG datasets which each consist of a large document collection with associated user requests. We focus on English texts, set $N_{e}=200$ and $N_{c}=50$, and use Llama3.3-70B-Instruct [18] as the generation and verification model to balance costs and quality. [^5] As we show in §5, creating effective CRUMQs does depend on proprietary models. This leads to a total of 3,048 CRUMQs. We next use UAEval4RAG to generate baseline out-of-database RAG queries over the same two datasets, leading to 7,559 out-of-database queries. Finally, we utilize the MultiHop-RAG dataset as-is, which consists of 2,556 multi-hop RAG queries and their associated gold contexts. Comparative dataset statistics as well as sample CRUMQs are provided in §5.
Unanswerability Evaluation. To verify the unanswerability value of our pipeline, we adopt the same experimental setup as in [29] and benchmark the same four leading RAG systems from [29] on the CRUMQs and UAEval4RAG queries, downsampled to 3,048 examples for fair comparison. We assume no access to external documents and use LlamaIndex [27] for RAG system implementation. We additionally follow [29] to evaluate RAG system performance when using leading proprietary LLMs Gemini-2.5-Pro [17], GPT-4o, and GPT-5 for generation.
Cheatability Evaluation. To demonstrate the robustness of our queries against reasoning shortcuts, we compare our CRUMQs (downsampled to 2,556 examples) against MultiHop-RAG [37]. Following [38], we first obtain LLM predictions for each dataset in the oracle setting on the tasks of answer prediction and paragraph-level support identification. We use Llama3.1-8B-Instruct, Llama3.3-70B-Instruct, GPT-4o, and Gemini-2.5-Pro to ensure coverage of diverse model types and sizes. We then repeat the same experiment on the DiRe probe [38] of each dataset, which is a dataset transformation designed to measure cheatability by gauging the extent to which correct answers can be achieved via reasoning over disjointed input contexts. Comparison of probe to non-probe performance serves as a model-agnostic estimate of cheatability; additional details are in [38].
@@ 5 Results and Analysis
Table 1: Performance of leading RAG systems [29] using different embedding, retrieval, reranking, and rewriting methods with GPT-4o generation on CRUMQs vs. UAEval4RAG queries. denotes Holm-Bonferroni corrected significance ($p<0.05$).
@@
Table 2: Performance of leading RAG systems [29] using different proprietary models for generation on CRUMQs vs. UAEval4RAG queries. As in [29], no reranking or rewriting is used. denotes Holm-Bonferroni corrected significance ($p<0.05$).
@@
Compared to prior multi-hop RAG benchmarks, CRUMQs are significantly more difficult and much less cheatable, with all comparisons significant after Holm–Bonferroni correction ($p<0.05$). As indicated by the black points in Fig. 3, we observe that CRUMQs pose a much greater challenge for leading LLMs than MultiHop-RAG across all tasks, with models achieving answer prediction scores up to only 31.9 versus up to 84.4 for MultiHop-RAG. Moreover, up to 96.7% of queries in MultiHop-RAG can be answered via disconnected reasoning, compared to only up to 23.6% of our CRUMQs. A similar trend is observed for the support identification task. These results confirm the efficacy of our pipeline for creation of challenging multi-hop RAG queries. [^6] We provide sample queries generated via the CRUMQs pipeline in Table 3 and comparative dataset statistics in Tables 4 and 5.
@@ 6.0.1 Acknowledgements
Table 3: Sample RAG queries generated via the CRUMQs pipeline.
# Hops Question
2 What historical event shares a name with a protocol used in computer science to achieve agreement among parties?
3 What is the impact of the estimated average annual loss of ice in Greenland and Antarctica combined on rising sea levels and ocean chemistry?
5 If the number of iterations in a stochastic linear bandit problem is equal to the number of major crusades, what would be the minimum parameter values that would allow for a ‘crusade’ against suboptimal solutions?
Table 4: Complexity and lexical diversity of CRUMQs versus prior RAG benchmarks.
Avg Gold Avg # Unique # # Verb-Noun
Ans Len Hops Verbs Uniq. Pairs
UAEval4RAG [29] 67.29 1.00 155 717
MultihopRAG [37] 1.32 2.38 109 332
CRUMQs (Ours) 20.04 2.74 225 826
Table 5: Comparative distribution of the number of hops per query.
1 2 3 4 5 6 7
MultiHop-RAG [37] 11.80% 42.20% 30.50% 15.60% 0.00% 0.00% 0.00%
CRUMQs (Ours) 11.50% 31.90% 31.60% 21.60% 3.10% 0.20% 0.10%